OpenStudy (anonymous):

What is the solution to the equation -3(h+5)+2=4(h+6)-9?

OpenStudy (anonymous):

\[ -3(h+5) + 2 = 4(h+6) - 9 \\ -3h -15 + 2 = 4h+24 - 9 \\ -3h -4h = 15 - 2 +24 - 9 \\ -7h = 28 \\ h = -\frac{28}{7} = -4 \]

OpenStudy (anonymous):

thank you @pitamar

OpenStudy (anonymous):

sure np